# Quantum Entrepreneurship Lab: Training a Future Workforce for the Quantum Industry

### v1

@@ I Introduction
Quantum technology is rapidly becoming a reality with the 2020s seeing significant investment from both the public and private sectors [1, 2, 3, 4]. From relatively mature technological solutions such as quantum-inspired computational techniques [5, 6] to the rapidly evolving quantum computing field [7, 8], a transformative industry is developing with disruptive innovation across science and engineering. While we can see this on the horizon, the technology is slow to mature and we do not necessarily know what form these industrial solutions will eventually take [9].
@@ II What institutions are involved?
II.1 Munich Quantum Valley
The Munich Quantum Valley (MQV) is an ecosystem intended to promote quantum technologies, connect researchers with industry, and ultimately, build a full-stack quantum computer in the German state of Bavaria [10]. This provides both a funding structure and cultivates a cooperative environment to connect quantum-related groups within Munich. The MQV is key for providing the network and connections between Munich’s universities, established industry with interest in quantum, as well as new quantum companies. The MQV is a collaboration between several institutions, in particular, the Technical University of Munich, Ludwig-Maximilians Universität München (University of Munich, LMU), Friedrich-Alexander University (FAU), Max Planck Institute of Quantum Optics (MPQ), Fraunhofer Institutes, and the German Aerospace Center (DLR) such that it provides access to a full quantum network. The QEL actively uses this collaborative network for funding, project ideas, as well as a channel for both incoming and outgoing quantum talent.
II.2 Technical University of Munich
@@
II.3 TUM Venture Lab Quantum / Semicon
@@ III.1 Objectives
III.1.1 Introducing students to research early
Early engagement in research is expected to produce more impactful outcomes in students’ future careers, whether they pursue academic research or industry roles [11, 12]. The QEL offers an opportunity for students to acquire and apply skills beyond traditional classroom settings by engaging with genuine research topics, thereby allowing them to explore their interests in a structured yet innovative environment.
III.1.2 Introducing students to deep tech
Deep tech [13] encompasses cutting-edge fields such as quantum computing, artificial intelligence, space technology, and biotechnology. These industries face significant challenges and offer transformative potential; however, the nascent state of these technologies often complicates the development of concrete ideas and business models. The QEL serves as a direct entry point into the deep tech landscape, acting as an essential stepping stone for translating innovative concepts into practical solutions.
III.1.3 Producing research output for academic groups
@@ III.1.4 Interdisciplinary collaborations
III.1.5 Industry involvement and mentorship
@@
III.1.6 Practical prototyping and proof-of-concepts
@@
III.1.7 Creating a quantum workforce
@@
III.2 Selecting industry-relevant topics
@@
III.2.1 Quantum-inspired techniques
This category focuses on advanced computational methods, such as tensor networks, machine learning algorithms, and optimization techniques, that originated from quantum science, often developed in parallel with the advent of quantum computers. Traditionally applied in physics, these techniques are now being explored in machine learning [14], fluid dynamics [6], generative AI [15], and data science [16] demonstrating their potential to drive innovation across multiple fields.
III.2.2 Solutions for the quantum industry
Rather than viewing the quantum industry solely as a collection of competing entities, we see it as a collaborative ecosystem working toward common objectives, such as achieving a stable quantum computer. Amid significant investment, there exists a substantial opportunity to develop products and services for quantum companies, for example, tailor-made simulators [17], compilers [18], or consulting on error mitigation and correction strategies [19].
By focusing on these two categories, the QEL ensures that student projects not only tackle real-world challenges but also contribute to the evolving landscape of the quantum industry. However, we expect that the topics in this course will fluctuate with the changing needs of the quantum landscape, in particular, including more hardware-focused topics, hybrid classical-quantum computing [20, 21], and applications of quantum computing [22]. In the end, the following topics were provided:
@@ III.3.3 Diversity and inclusion
Figure 2: Overall course timeline described in Sec. III.4
III.4 Structure
@@
III.4.1 Kick-off phase (2 weeks)
@@
III.4.2 Ideation phase (2 months)
@@
III.4.3 Development phase (1 month)
@@
III.4.4 Delivery phase (2 months)
@@ IV Outcomes
IV.1 Quantum-inspired data science
This project explored the use of tensor networks for data compression, addressing the increasing computational and energy demands of artificial intelligence. The team developed a quantum-inspired approach using matrix product states (MPS) [23] to compress the required storage space of data sets, with potential applications in optimizing machine learning models, particularly large-scale neural networks.
@@ IV.1 Quantum-inspired data science
IV.2 Quantum compiler development
Students in this project focused on optimizing quantum compilation for superconducting qubits [24], with the goal of improving algorithm performance on noisy intermediate-scale quantum (NISQ) devices [7]. Their approach aimed to reduce computational overhead and enhance resilience against noise through specialized compiler techniques.
@@ IV.2 Quantum compiler development
IV.3 Machine learning for quantum chemistry
This project applied machine learning to accelerate quantum chemistry simulations, particularly in the context of materials discovery. The students developed models to streamline key computational processes, such as band structure calculations and Green’s function predictions, which are crucial in material science research [25, 26].
@@ IV.3 Machine learning for quantum chemistry
IV.4 Quantum-inspired generative AI
Leveraging techniques from quantum many-body physics, this project explored the application of matrix product states (MPS) [23] to generative AI models, with a focus on financial applications [27]. The students investigated how these models could improve predictive capabilities in areas such as stock and option pricing.
@@ IV.4 Quantum-inspired generative AI
IV.5 Quantum hardware/error correction codesign
This project examined quantum error correction (QEC) strategies tailored to different hardware architectures [28]. The students developed methods to adapt QEC techniques based on the specific characteristics of various quantum platforms, such as superconducting qubits [24], trapped ions [29], and neutral atoms [30].
@@ IV.5 Quantum hardware/error correction codesign
IV.6 Student achievements and impact
@@
IV.7 Collaboration with industry and research institutions
@@ VI Future improvements
VI.1 Re-integrating industry partners
@@
VI.2 Improving business education
@@
VI.3 Expanding networking and mentorships
@@ Acknowledgments
In particular, we would like to acknowledge the participants of the course including Egor Agapov, Oriol Bertomeu, Andrés Carballo, Rhea Daga, Raunaq Jain, Danylo Kolesnyk, Jessica Link, Alexander Orlov, Daniel Ortmann, Mert Özel, Alan Piovesana, Max Gert Richter, Leon Seidelmeier, Agata Skoczylas, Laurin Steiner, Andreas Strobl, Myron Theocharakis, Dimitrios Vasileiadis, and Yelyzaveta Vodovozova.

### v2

@@ Abstract
Index Terms:
Quantum Technology, Entrepreneurship, Deep Tech, Education
@@ I Introduction
Quantum technology is rapidly becoming a reality with the 2020s seeing significant investment from both the public and private sectors [jurczak_investing_2023, mckinsey2024quantum, putranto_deep_2024, sotelo_quantum_2021]. From relatively mature technological solutions such as quantum-inspired computational techniques [altmann_quantum-inspired_2018, peddinti_quantum-inspired_2024] to the rapidly evolving quantum computing field [preskill_quantum_2018, deshpande_assessing_2022], a transformative industry is developing with disruptive innovation across science and engineering. While we can see this on the horizon, the technology is slow to mature and we do not necessarily know what form these industrial solutions will eventually take [martens_acceptance_2025].
@@ II What institutions are involved?
II-A Munich Quantum Valley
The Munich Quantum Valley (MQV) is an ecosystem intended to promote quantum technologies, connect researchers with industry, and ultimately, build a full-stack quantum computer in the German state of Bavaria [MQV2024strategy]. This provides both a funding structure and cultivates a cooperative environment to connect quantum-related groups within Munich. The MQV is key for providing the network and connections between Munich’s universities, established industry with interest in quantum, as well as new quantum companies. The MQV is a collaboration between several institutions, in particular, the Technical University of Munich, Ludwig-Maximilians Universität München (University of Munich, LMU), Friedrich-Alexander University (FAU), Max Planck Institute of Quantum Optics (MPQ), Fraunhofer Institutes, and the German Aerospace Center (DLR) such that it provides access to a full quantum network. The QEL actively uses this collaborative network for funding, project ideas, as well as a channel for both incoming and outgoing quantum talent.
II-B Technical University of Munich
@@
II-C TUM Venture Lab Quantum / Semicon
@@ III-A Objectives
III-A1 Introducing students to research early
Early engagement in research is expected to produce more impactful outcomes in students’ future careers, whether they pursue academic research or industry roles [rodenbusch2016, bowman_getting_2018]. The QEL offers an opportunity for students to acquire and apply skills beyond traditional classroom settings by engaging with genuine research topics, thereby allowing them to explore their interests in a structured yet innovative environment.
III-A2 Introducing students to deep tech
Deep tech [mit2023deeptech] encompasses cutting-edge fields such as quantum computing, artificial intelligence, space technology, and biotechnology. These industries face significant challenges and offer transformative potential; however, the nascent state of these technologies often complicates the development of concrete ideas and business models. The QEL serves as a direct entry point into the deep tech landscape, acting as an essential stepping stone for translating innovative concepts into practical solutions.
III-A3 Producing research output for academic groups
@@ III-A4 Interdisciplinary Collaborations
III-A5 Industry Involvement and Mentorship
@@
III-A6 Practical Prototyping and Proof of Concepts
@@
III-A7 Creating a quantum workforce
@@
III-B Selecting industry-relevant topics
@@
III-B1 Quantum-inspired techniques
This category focuses on advanced computational methods, such as tensor networks, machine learning algorithms, and optimization techniques, that originated from quantum science, often developed in parallel with the advent of quantum computers. Traditionally applied in physics, these techniques are now being explored in machine learning [huynh_quantum-inspired_2023], fluid dynamics [peddinti_quantum-inspired_2024], generative AI [moussa_application_2023], and data science [cichocki_tensor_2014] demonstrating their potential to drive innovation across multiple fields.
III-B2 Solutions for the quantum industry
Rather than viewing the quantum industry solely as a collection of competing entities, we see it as a collaborative ecosystem working toward common objectives, such as achieving a stable quantum computer. Amid significant investment, there exists a substantial opportunity to develop products and services for quantum companies, for example, tailor-made simulators [xu_herculean_2023], compilers [maronese_quantum_2021], or consulting on error mitigation and correction strategies [filippov_scalability_2024].
By focusing on these two categories, the QEL ensures that student projects not only tackle real-world challenges but also contribute to the evolving landscape of the quantum industry. However, we expect that the topics in this course will fluctuate with the changing needs of the quantum landscape, in particular, including more hardware-focused topics, hybrid classical-quantum computing [yang_semiconductor_2022, ovalle_magallanes_hybrid_2022], and applications of quantum computing [bova_commercial_2021]. In the end, the following topics were provided:
@@ III-C3 Diversity and Inclusion
Figure 2: Overall course timeline described in Sec. III-D
III-D Structure
@@
III-D1 Kick-off phase (2 weeks)
@@
III-D2 Ideation phase (2 months)
@@
III-D3 Development phase (1 month)
@@
III-D4 Delivery phase (2 months)
@@ IV Outcomes
IV-A Quantum-inspired data science
This project explored the use of tensor networks for data compression, addressing the increasing computational and energy demands of artificial intelligence. The team developed a quantum-inspired approach using matrix product states (MPS) [orus_practical_2014] to compress the required storage space of data sets, with potential applications in optimizing machine learning models, particularly large-scale neural networks.
@@ IV-A Quantum-inspired data science
IV-B Quantum compiler development
Students in this project focused on optimizing quantum compilation for superconducting qubits [arute_quantum_2019], with the goal of improving algorithm performance on noisy intermediate-scale quantum (NISQ) devices [preskill_quantum_2018]. Their approach aimed to reduce computational overhead and enhance resilience against noise through specialized compiler techniques.
@@ IV-B Quantum compiler development
IV-C Machine learning for quantum chemistry
This project applied machine learning to accelerate quantum chemistry simulations, particularly in the context of materials discovery. The students developed models to streamline key computational processes, such as band structure calculations and Green’s function predictions, which are crucial in material science research [kuramoto_quantum_2020, sheridan_data-driven_2021].
@@ IV-C Machine learning for quantum chemistry
IV-D Quantum-inspired generative AI
Leveraging techniques from quantum many-body physics, this project explored the application of matrix product states (MPS) [orus_practical_2014] to generative AI models, with a focus on financial applications [mugel_use_2020]. The students investigated how these models could improve predictive capabilities in areas such as stock and option pricing.
@@ IV-D Quantum-inspired generative AI
IV-E Quantum hardware/error correction codesign
This project examined quantum error correction (QEC) strategies tailored to different hardware architectures [roffe_quantum_2019]. The students developed methods to adapt QEC techniques based on the specific characteristics of various quantum platforms, such as superconducting qubits [arute_quantum_2019], trapped ions [cirac_quantum_1995], and neutral atoms [wintersperger_neutral_2023].
@@ IV-E Quantum hardware/error correction codesign
IV-F Student achievements and impact
@@
IV-G Collaboration with Industry and Research Institutions
@@ VI Future Improvements
VI-A Re-integrating industry partners
@@
VI-B Improving business education
@@
VI-C Expanding Networking and Mentorships
@@ Acknowledgments
We would like to thank Raúl Berganza Gómez and Tom Hubregtsen for laying the foundation of the original Quantum Entrepreneurship Lab, as well as Benjamin Schiffer for his insights.
We would also like to acknowledge the participants of the course including Egor Agapov, Oriol Bertomeu, Andrés Carballo, Rhea Daga, Raunaq Jain, Danylo Kolesnyk, Jessica Link, Alexander Orlov, Daniel Ortmann, Mert Özel, Alan Piovesana, Max Gert Richter, Leon Seidelmeier, Agata Skoczylas, Laurin Steiner, Andreas Strobl, Myron Theocharakis, Dimitrios Vasileiadis, and Yelyzaveta Vodovozova.
@@
[…]